Use Strassen’s algorithm to compute the matrix product

Show your work.
Answer #1

`Hey,

Note: If you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

The first matrices are Si = 6 S2 = 4 S6 = 8 S7 = -2 Sg = 6 Sy = -6 S10 = 14. = -2 S5 = 6 The products are R =1.6 = 6 P = 4.2

The four matrices are Cu = 48 + (-10-8+(-12) = 18 C12 = 6+8 = 14 C21 = 72+(-10) = 62 C22 = 48 + 6 – 72-(-84) = 66. The result

Kindly revert for any queries

Thanks.